OpenStudy (anonymous):

The graph of y = -3x2- 2x - 5 has a: A. Maximum B. Minimum Find the vertex of the graph of the following equation: y = x2 + 8x -12. A. (4, -28) B. (-4, -28) C. (-4, 4) D. (4, 4)

OpenStudy (the_fizicx99):

For A, plot it. If it has a n shaped curve it has a maximum, if it has a U curve it has a minimum. Yours has a maximum. as for #2, you can use the formula \(\ \sf -(\dfrac{b}{2a}) \) In \(\ \sf \Large y = x^2 + 8x - 12 \) \(\ \sf \Large B \rightarrow 8 \) \(\ \sf \Large A \rightarrow 1 \) \(\ \sf \Large -(\dfrac{8}{2(1)}) = -\dfrac{8}{2} = -4\) Now plug in \(\ \sf \color{green}{-4}\) into \(\ \sf \Large f(\color{green}{x}) = \color{green}{x}^2 + 8\color{green}{x} - 12 \) and that'll be your coordinates to the vertex.

OpenStudy (the_fizicx99):

The formula is actually \(\ \sf \color{green}{x} = -(\dfrac{b}{2a}) \)

OpenStudy (anonymous):

\[F(-4)-4^{2}+8(-4)-12\] @tHe_FiZiCx99

OpenStudy (the_fizicx99):

What's that?

OpenStudy (anonymous):

i pluged in -4 into the green spaces

OpenStudy (the_fizicx99):

Yeah I thought that might help :> Yeah you're right now simplify it, don't forget the equal sign! f(-4) = (-4)^2+ 8(-4) - 12

OpenStudy (the_fizicx99):

What did you get?

OpenStudy (anonymous):

i got C @tHe_FiZiCx99

OpenStudy (the_fizicx99):

No no, what did you get when you plugged in -4 into the function?

OpenStudy (anonymous):

f(-4) = (-4)^2+ 8(-4) - 12 -4x^2+16-12

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Um no xD What is -4^2 ? Btw, I think what confused you is the x, -4 takes the x place because x IS -4

OpenStudy (anonymous):

-4*-4=16

OpenStudy (the_fizicx99):

Yes, now what is 8(-4) = ?

OpenStudy (anonymous):

32

OpenStudy (the_fizicx99):

Nope, negative sign!

OpenStudy (anonymous):

-32

OpenStudy (the_fizicx99):

Yes, now: 16 - 32 - 12 = ?

OpenStudy (anonymous):

-28

OpenStudy (the_fizicx99):

Yes! Now the cool part in this, remember the formula \(\ \sf x = -(\dfrac{b}{2a}) \) ? Well when you solve that it gives you x = -4, now when you plugged that into the function, you got -28 So the coordinates are (-4,-28), this is where the vertex is!

OpenStudy (anonymous):

thanksso much for the help

OpenStudy (the_fizicx99):

You're welcome :>